OpenStudy (anonymous):

Find the volume of the circular cylinder. Radius = 2 & Height = 12 .

Parth (parthkohli):

The volume of a cube is \(\pi r^2h\)

OpenStudy (anonymous):

Remember what I said before? You can just find the area and then multiply by the height. Do you remember the formula for the area of a circle?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \pi \cdot 2 \cdot 2 \cdot 12 }\)

OpenStudy (anonymous):